OpenStudy (xxmuslimxx):

Can You Check My Work? Find the average for the following numbers: A) 12 11 1/2 11 2/8 = 12 1/2 B) 11 7 9 = 9 C) 12 13 10 = 11 67/100

OpenStudy (anonymous):

I know that B is correct.

OpenStudy (xxmuslimxx):

Yeah....Uh...I already knew that one, I was looking forward for help with A and C...Thanks anyway Stace.

OpenStudy (texaschic101):

A. (12 + 11 1/2 + 11 2/8) / 3 = (12 + 11.5 + 11.25) / 3 = 34.75/3 = 11.58 which is the same as 11 58/100 which equals 11 29/50 C. (12 + 13 + 10) / 3 = 35/3 = 11 2/3

OpenStudy (xxmuslimxx):

Thank you so much! That was really helpful.